• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jwexler

    Log entry time 07:59:22 on June13,2010

    Entry number 326259

    keyword=Shift Summary - Owl 6/13/2010

    Shift Summary: Owl 6/13/2010

    00:00 - Putting in PbD#1. Will do position scan then transverse asymmetry production.

    1:23 - There was a problem with the Hall A line, MCC fixed this by changing the optics. This also probably changed the magnet settings, so we will do another position scan

    3:00 - Positions OK, raster on, cooler on, heater off. Going to ramp up *very slowly* (1/3rd the ramp rate) to 10uA, see how vacuum fares.

    4:00 - Ramped to 40uA, started a run. Saw signals were low, changed HVs. Lost control of right arm HVs again (telnet still in 'packed' mode, HV GUI not starting). Will try to solve this, but will continue slow ramp up to 50uA.

    4:20 - Still struggling with HVs. Will run at 50uA using "100k channels at 40uA" settings. Note the lumis are saturating.

    4:30 - HV GUI started, changing HVs for lumis/detectors for nominal 50uA, 100k channel running.

    5:00 - Det. HVs changed (HV GUI difficulties slowed that down considerably). Started transverse asymmetry production running on PbD#1.

    6:40 - MCC called to say that the temperature of the Hall A bcm has been slightly wrong since 6:00, and therefore the current delivered to the hall might be slightly different than 50uA. They have called an expert to look at it. Of course, we will continue to take beam.

    7:30 - Changed slugs to IHWP IN (from OUT).

    7:48 - Target vacuum excursion. It went up to 9e-5. We requested for MCC to take away beam and we will ramp up slowly after the vacuum settles.